-6x+16=5(x+1)
Please tell me how to solve this problem.
Thank you

Answer by jcamille(5)

-6x+16=5(x+1)
-6x+16=5x+5
-6x+16-5x-16=5x+5-5x-16
combine like terms
-6x-5x+16-16=5x+5-16
-11x=-11
-11x/-11=-11/-11
x=1